Calculate the final temperature, in degrees Celsius, ∘C, for each of the following, if V and n do not change:

You may want to reference (Pages 286 - 288) Section 8.4 while completing this problem.

1.

A sample of helium gas with a pressure of 330 torr at 0 ∘C is heated to give a pressure of 2100 torr.

Express your answer to two significant figures and include the appropriate units

t=

2.

A sample of air at 13 ∘C and 750. mmHg is cooled to give a pressure of 660. mmHg.

Express your answer as integer and include the appropriate units.

t=
Answer #1

1)

Given:

Ti = 0.0 oC

= (0.0+273) K

= 273 K

Pi = 330 torr

Pf = 2100 torr

use:

Pi/Ti = Pf/Tf

330.0 torr / 273.0 K = 2100.0 torr / Tf

Tf = 1737 K

Answer: 1.7*10^3 K

2)

Given:

Ti = 13.0 oC

= (13.0+273) K

= 286 K

Pi = 750 mmHg

Pf = 660 mmHg

use:

Pi/Ti = Pf/Tf

750.0 mmHg / 286.0 K = 660.0 mmHg / Tf

Tf = 252 K

Answer: 252 K